OpenStudy (anonymous):

Assume that the family income in a certain town is distributed normally with mean $60000 and a standard deviation of $10000. If the poverty level is $30000, what percentage of the family live above the poverty line? Recall that F(z)=(1/(2pi)^1/2)integral(e^(-x^2/2))dt from z to -infinity

OpenStudy (anonymous):

Let \(X\) be the variable to represent the family income. You want to find \(P(X>30,000)\). First transform to a standard normal distribution: \[P(X>30,000)=P\left(\frac{X-60,000}{10,000}>\frac{30,000-60,000}{10,000}\right)=P(Z>-3)\]

OpenStudy (anonymous):

the answer choices are a) F(5/3) b)F(-3) c)1-F(3) d)1-F(-5/3) e)F(3)

OpenStudy (anonymous):

Then \[\large \begin{align*}P(Z>-3)&=1-P(Z<-3)\\&=1-F(-3)\\ &=1-\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{-3}\exp\left(-\dfrac{t^2}{2}\right)~dt \end{align*}\] If I remember correctly, \(P(Z>-3)\) is very close to 1 as a probability...

OpenStudy (anonymous):

Ah, they want you to use the symmetry of the standard normal distribution here. \(F(3)=1-F(-3)\).

OpenStudy (anonymous):

Oh my god thank you so much!!!

OpenStudy (anonymous):

yw